OpenStudy (anonymous):

Trying to figure out if im doing this right? Let f(x)=-5x+3 and g(x)=x^2-2x+1 f(-3)= g(-x)= f(4m-2)= g(x+2)= Please Help me understand

Nnesha (nnesha):

\[\large\rm f(\color{reD}{-3}) \rightarrow \color{reD}{x=-3}\] f(-3) means x = -3 so just substitute x for -3 into the f(x) function ssame steps for other 3 :=))]

OpenStudy (anonymous):

So I got 18 for the first one but the others are confusing my professor just gave it to us and is due in the am

Nnesha (nnesha):

huh ohhh professor :P alright like i say same steps for g(-x) :=)) what you should plug into the g(x) for x ?

Nnesha (nnesha):

\[\large\rm f(\color{reD}{-x}) \rightarrow \color{reD}{x=??}\]

OpenStudy (anonymous):

See I don't. Understand it would just be a negative x right no number makes it a little harder

Nnesha (nnesha):

ye just replace `x` with `-x` \[\rm g(\color{reD}{-x}) =\color{Red}{x}^2-2\color{reD}{x}+1\]

Nnesha (nnesha):

remember when we take even power of a negative number/variable u will get positive answer :=))

Nnesha (nnesha):

to find g(x+2) substitute x for x+2 \[\large\rm f(\color{reD}{x+2}) \rightarrow \color{reD}{x=(x+2)}\] \[\huge\rm g(\color{red}{x+2})= \color{ReD}{x}^2-2\color{ReD}{x}+1\] \[\large\rm g(\color{red}{x+2})= \color{ReD}{(x+2)}^2-2\color{ReD}{(x+2)}+1\] for this one (x+2)^2 is same as (x+2)(x+2 )multiply them using foil method and distribute (x+2) by -2 at the end just combine like terms same steps for number 3 try it!!! :=))

OpenStudy (anonymous):

Whats the answer for f(4m-2)

Nnesha (nnesha):

what did you get ?? f(4m-2) is easier thang(x+2) :=)) substitute x for `4m-2` into f(x) function

OpenStudy (anonymous):

3 what's the answer